Calculate the value of the ionization constant Kb if the pKb is 8.67


1
Expert's answer
2021-08-23T02:57:14-0400

Solution:

pKb = - log(Kb)

Kb = 10-pKb

Kb = 10-8.67 = 2.13×10-9

Kb = 2.13×10-9


Answer: The value of the ionization constant (Kb) is 2.13×10-9